OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2). A. y = 3x - 5 B. y = 3x - 2 C. y = 3x + 1 D. y = 3x + 3

OpenStudy (anonymous):

-2 = 3(1) + b -2 = 3 + b, so b = -5 The line equation is y = 3x - 5. So it was A. Hope this helps!